# Longitudinal efficient adjustment sets for time-varying treatment effect estimation in nonparametric causal graphical models

### v1

@@ Abstract
Criteria for identifying optimal adjustment sets (i.e., yielding a consistent estimator with minimal asymptotic variance) for estimating average treatment effects in parametric and nonparametric models have recently been established. In a single treatment time point setting, it has been shown that the optimal adjustment set can be identified based on a causal directed acyclic graph alone. In a longitudinal treatment setting, previous work has established graphical rules to compare the asymptotic variance of estimators based on nested time-dependent adjustment sets. However, these rules do not always permit the identification of an optimal time-dependent adjustment set based on a causal graph alone. In this paper, we extend previous results by exploiting conditional independencies that can be read from the graph. We demonstrate theoretically and empirically that our results can yield estimators with even lower asymptotic variance than those allowed by previous results. We conjecture that our new results may even allow the identification of an optimal time-dependent adjustment set based on the causal graph and provide numerical examples supporting this conjecture.
Keywords adjustment sets  $\cdot$ back-door formula  $\cdot$ causal inference  $\cdot$ nonparametric inference
@@ 1 Introduction
One of the most important challenges for causal inference when conducting observational studies lies in the ability to control for confounding. In practice, investigators often identify confounding covariates during study design and estimate the causal effect of treatment or exposure by controlling, or adjusting, for a subset of covariates using an appropriate adjustment method during data analysis. One strategy for selecting which variables to control first entails constructing a causal graph based on expert opinion and subject-matter knowledge. Then, the backdoor criterion can be used to identify sets of variables that, when adequately controlled, allow for the elimination of confounding bias [1]. Adjustment sets that allow consistent estimation of a causal effect are known as “sufficient adjustment sets”.
@@
For a given causal graph, several sufficient adjustment sets may exist. A popular strategy for choosing among them is to select a minimal sufficient adjustment set, that is, the sufficient adjustment set that includes the fewest variables [2, 3]. Such a strategy can be appealing to proponents of parsimonious models but offers no guarantee that the resulting treatment effect estimator has a lower variance than an estimator adjusting for more variables. Alternatively, recent work has focused on the identification of the sufficient adjustment set that results in a treatment effect estimator with the lowest possible variance. Notably, Henckel et al[4] developed a criterion to identify the optimal sufficient adjustment set for estimating the effect of a point exposure on an outcome. This criterion is based on a causal graph and assumes that the relations between the variables in the graph are linear. Rotnitzky and Smucler [5] extended those results in several directions, for instance by showing that the criterion proposed by Henckel et al[4] is valid for a broad class of nonparametric estimators of the average treatment effect. They further provided a new graphical criterion for comparing adjustment sets in terms of estimation variance for estimating the effect of a time-dependent treatment subject to time-dependent confounding [5].
@@
In this paper, we extend previous results by proposing an alternative definition of a sufficient time-dependent adjustment set and a novel criterion for comparing the variance of estimators based on different time-dependent adjustment sets. We show that employing our criterion can further decrease the variance of estimators of time-dependent treatment effects. The rest of the paper is organized as follows. In Section 2 we review some background concepts, and present the assumed data structure and the notation. In Section 3 we provide a novel definition of time-dependent adjustment sets. In Section 4 we present our new criterion and provide numerical illustrations comparing the variance of estimators based on our proposed criterion and those based on previous results. Section 5 ends with a conclusion. All proofs of results stated in the main text can be found in the Appendix.
@@ 2 Background
We begin by briefly reviewing some concepts of causal graph theory and introducing some notation. The unfamiliar reader is referred to other references for a more detailed introduction (e.g., the appendix of [6], or [1]). Let $\mathcal{G}$ be a directed acyclic graph (DAG) with nodes $\mathbf{V}$ representing the variables, that is, a graph for which all edges are unidirectional arrows and for which there is no cycle. A path is defined as a sequence of variables connected by arrows. A variable is said to be a collider in a given path if two arrows point to this variable in this path. Denote by $P$ the unknown distribution of the variables represented in the graph and by $Pa_{\mathcal{G}}(v_{k})$ the parents of $v_{k}$ according to the graph $\mathcal{G}$, that is, $Pa_{\mathcal{G}}(v_{k})$ is the set of variables in $\mathcal{G}$ that have arrows stemming from them and pointing into $v_{k}$. We say that a DAG is a causal graph if all common causes of the variables depicted in the graph are also on the graph [1] and if
$P=\prod_{v_{k}\in\mathbf{V}}f\left(v_{k}\mid Pa_{\mathcal{G}}(v_{k})\right),$
@@
Suppose that we have a study with $p+1$ follow-up times and $n$ individuals sampled from a population so that the time-ordered (or otherwise topologically ordered) data take the following form $\mathcal{O}=(\mathbf{C}_{0},A_{0},\mathbf{C}_{1},A_{1},...,\mathbf{C}_{p},A_{p},Y)$ where $\mathbf{A}=(A_{0},...,A_{p})\subset\mathbf{V}$ is a time-varying exposure or treatment, $\mathbf{C}_{k}=(C_{k,1},C_{k,2},...,C_{k,n_{ck}}),k=0,...,p$ is a vector of time-dependent covariates (with $n_{ck}$ the number of covariates at time $k$) and $Y\subset\mathbf{V}$ is the outcome of interest, measured at the end of the study. We assume that $A_{k},k=0,...,p$, correspond to finite valued random variables. For time ordered sets of variables, we use an overbar to denote the history of that variable up to a given time point. For example $\bar{\mathbf{A}}_{p}=(A_{0},...,A_{p})$. Let $Y^{\mathbf{a}}$ be the counterfactual outcome, that is, the outcome that would have been observed if, possibly contrary to fact, exposure had been $\mathbf{A}=\mathbf{a}$ where $\mathbf{a}=(a_{0},...,a_{p})$.
Based on a causal graph, it is possible to construct a single world intervention graph (SWIG) that represents the causal model (including counterfactual variables) under a hypothetical intervention [7, 8]. As will be seen shortly, we will be using these SWIGs to identify sets of variables that meet the sequential conditional exchangeability assumption of causal effects. A SWIG can be constructed from a DAG by splitting each treatment node $A_{0},...,A_{p}$ in two parts, one part for the observed and one for the value set by the (hypothetical) intervention: $A_{0}\mid a_{0},...,A_{p}\mid a_{p}$. All arrows entering into a node $A_{k}$ in the DAG also enter $A_{k}$ in the SWIG, whereas all arrows stemming from a node $A_{k}$ in the DAG now stem from $a_{k}$. Further, all descendants of a node $a_{k}$ in the SWIG are indexed by an $a_{k}$ superscript to emphasize that it is a counterfactual variable. Figure 1 presents an example of a DAG and its associated SWIG.
[…]
@@
Independence between variables can be read from a causal DAG or a SWIG using the concept of d-separation [1]. Consider disjoint nodes, $V_{j}$ and $V_{k}$, and set of nodes $\mathbf{V}_{\ell}$. The nodes $V_{j}$ and $V_{k}$ are d-separated by $\mathbf{V}_{\ell}$ if all paths between $V_{j}$ and $V_{k}$ are blocked given $\mathbf{V}_{\ell}$. A path is said to be blocked given $\mathbf{V}_{\ell}$ if a) it contains a non-collider that is an element of $\mathbf{V}_{\ell}$, or b) there is a collider on the path where neither it nor its descendants are in $\mathbf{V}_{\ell}$. If $V_{j}$ and $V_{k}$ are d-separated given $\mathbf{V}_{\ell}$, then $V_{j}$ and $V_{k}$ are also independent conditional on $\mathbf{V}_{\ell}$.
Our focus is on the efficient estimation of contrasts between counterfactual outcome expectations of the form $\Delta(P;\mathcal{G})\equiv\sum_{\mathbf{a}\in\mathcal{A}}c_{\mathbf{a}}\mathbb{E}_{P}(Y^{\mathbf{a}})$, where $\mathbf{a}$ is a vector taking values in the finite set $\mathcal{A}$ of all possible values that $\mathbf{A}$ can take and $c_{\mathbf{a}}\in\mathbb{R}$. For example in the context of a two time point study with a binary treatment, $\mathbf{A}=(A_{0},A_{1})$ and $\mathcal{A}=\{(0,0),(1,1),(1,0),(0,1)\}$. If $c_{(0,0)}=-1$, $c_{(1,1)}=1$, and $c_{(1,0)}=c_{(0,1)}=0$ then $\Delta(P;\mathcal{G})=\mathbb{E}_{P}(Y^{(1,1)})-\mathbb{E}_{P}(Y^{(0,0)})$. These causal contrasts can be identified from the observed data by first identifying the counterfactual expectations $\chi_{a}(P;\mathcal{G})\equiv\mathbb{E}_{P}(Y^{\mathbf{a}})$.
@@
Let ${\mathbf{Z}=(\mathbf{Z}_{0},...,\mathbf{Z}_{p})\subset\mathbf{C}}$ be time ordered, disjoint sets of variables. Similar to Rotnitzky and Smucler[5], we say that $\mathbf{Z}$ is a sufficient time-dependent adjustment set if
$\mathbb{E}_{P}(Y^{\mathbf{a}})$ $=\mathbb{E}_{P}(...\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid\bar{\mathbf{a}}_{p},\bar{\mathbf{Z}}_{p})\mid\bar{\mathbf{a}}_{p-1},\mathbf{\bar{Z}}_{p-1}\}...\mid a_{0},\mathbf{Z}_{0}])$ (1)
@@
$\mathbb{E}_{P}(Y^{\mathbf{a}})$ $=\mathbb{E}_{P}\left\{\frac{I_{\bar{\mathbf{a}}}(\bar{\mathbf{A}})Y}{\prod_{k=0}^{p}\mathbb{P}(A_{k}=a_{k}\mid\bar{\mathbf{a}}_{k-1},\bar{\mathbf{Z}}_{k})}\right\},$
@@
One way to identify a sufficient time-dependent adjustment set from a causal graph is to first construct a SWIG from the causal graph, then identify a set $\mathbf{Z}_{0},...,\mathbf{Z}_{p}$ such that the sequential conditional exchangeability assumption, ${Y^{\mathbf{a}}\perp\!\!\!\perp A_{k}\mid\bar{\mathbf{A}}_{k-1},\bar{\mathbf{Z}}_{k}}$ for $k=0,...,p$, holds, where $\bar{A}_{-1}\equiv\varnothing$. Example 1 illustrates this process.
@@
Assume that Figure 1 represents a causal DAG and its associated SWIG. Using d-separation rules, it can be seen that $Y^{a}\perp\!\!\!\perp A_{0}\mid C_{0,2}$ and that $Y^{a}\perp\!\!\!\perp A_{1}^{a_{0}}\mid C_{1,2}^{a_{0}},C_{0,2},A_{0}$, for example. Under consistency (i.e., if $A_{k}=a_{k}$ then $V^{a_{k}}=V$), the second conditional independence becomes ${Y^{a}\perp\!\!\!\perp A_{1}\mid C_{1,2},C_{0,2},A_{0}}$. As such, $\mathbb{E}_{P}(Y^{\mathbf{a}})=\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid A_{1}=a_{1},C_{1,2},A_{0}=a_{0},C_{0,2})\mid A_{0}=a_{0},C_{0,2}\}]$ and $\mathbf{Z}=(\mathbf{Z_{0}}=C_{0,2},\mathbf{Z_{1}}=C_{1,2})$ is a time-dependent adjustment set.
@@ 3 Proposal
Let $\mathbf{Z}=(\mathbf{Z}_{0},...,\mathbf{Z}_{p})\subset\mathbf{V}\backslash Y$. We say that $\mathbf{Z}$ is a sufficient time-dependent adjustment set if
$\mathbb{E}_{P}(Y^{\mathbf{a}})$ $=\mathbb{E}_{P}(...\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid a_{p},\mathbf{Z}_{p}(\bar{\mathbf{a}}_{p-1}))\mid a_{p-1},\mathbf{Z}_{p-1}(\bar{\mathbf{a}}_{p-2})\}...\mid a_{0},\mathbf{Z}_{0}])$ (2)
@@
$\mathbb{E}_{P}(Y^{\mathbf{a}})$ $=\mathbb{E}_{P}\left\{\frac{I_{\bar{\mathbf{a}}}(\bar{\mathbf{A}})Y}{\prod_{k=0}^{p}\mathbb{P}(A_{k}=a_{k}\mid\mathbf{Z}_{k}(\bar{\mathbf{a}}_{k-1}))}\right\}$
@@
where we use the notation $\mathbf{Z}_{k}(\bar{\mathbf{a}}_{k-1})$ to emphasize that if some or all past treatments are in $\mathbf{Z}_{k}(\bar{\mathbf{a}}_{k-1})$ then they are held fixed.
@@
Definition 2 differs from Definition 1 in a few subtle but important aspects. First, unlike Definition 1, Definition 2 allows for previous treatments to be members of $\mathbf{Z}_{k}$ (i.e., $\bar{A}_{k-1}\in\bar{\mathbf{Z}}_{k}$ is allowed), $k=1,...,p$. In addition, the expectations in the identification formula (2) are only conditional on $A_{k}$ and $\mathbf{Z}_{k}$, not on $\bar{A}_{k}$ and $\bar{\mathbf{Z}}_{k}$, $k=0,...,p$, as in Definition 1.

### v2

# Efficient adjustment sets for time-dependent treatment effect estimation in nonparametric causal graphical models

@@ Abstract
Criteria for identifying optimal adjustment sets yielding consistent estimation with minimal asymptotic variance of average treatment effects in parametric and nonparametric models have recently been established. In a single treatment time point setting, it has been shown that the optimal adjustment set can be identified based on a causal directed acyclic graph alone. In a time-dependent treatment setting, previous work has established graphical rules to compare the asymptotic variance of estimators based on nested time-dependent adjustment sets. However, these rules do not always permit the identification of an optimal time-dependent adjustment set based on a causal graph alone. We extend those results by exploiting conditional independencies that can be read from the graph and demonstrate theoretically and empirically that our results can yield estimators with lower asymptotic variance than those allowed by previous results. We further show how our results allow for the identification of optimal adjustment sets based on a directed acyclic graph alone in the time-dependent treatment setting.
Keywords adjustment sets $\cdot$ causal inference $\cdot$ nonparametric inference
@@ 1 Introduction
One of the most important challenges for causal inference with observational data lies in the ability to control for confounding. In practice, investigators often identify and measure confounding covariates during study design and estimate the causal effect of treatment or exposure by controlling, or adjusting, for a subset of covariates using an appropriate adjustment method during data analysis. One strategy for selecting which variables to control first entails constructing a causal graph based on expert opinion and subject-matter knowledge. Then, the backdoor criterion can be used to identify sets of variables that, when adequately controlled, allow for the elimination of confounding bias [1]. Adjustment sets that allow for consistent estimation of a causal effect are known as “sufficient adjustment sets”.
@@
For a given causal graph, several sufficient adjustment sets may exist. A popular strategy for choosing among them is to select a minimal sufficient adjustment set, that is, the sufficient adjustment set that includes the fewest variables [2, 3]. Such a strategy can be appealing to proponents of parsimonious models but offers no guarantee that the resulting treatment effect estimator has a lower variance than an estimator adjusting for more variables. Alternatively, recent work has focused on the identification of the sufficient adjustment set that results in a treatment effect estimator with the lowest possible asymptotic variance. Notably, Henckel et al [4] developed a criterion to identify the optimal sufficient adjustment set for estimating the effect of a single or joint exposure intervention on one or multiple outcomes. This criterion is based on a causal graph and assumes that the relations between the variables in the graph are linear. Rotnitzky and Smucler [5] extended those results in several directions, for instance by showing that the criterion proposed by Henckel et al [4] is valid for a broad class of nonparametric estimators and without assuming linear relations between variables for estimating the average effect of an exposure on an outcome. They further provided a new graphical criterion for comparing adjustment sets in terms of the variance of estimators of the effect of a time-dependent treatment subject to time-dependent confounding [5]. An important limitation of the latter result is the apparent impossibility to identify an optimal adjustment set based on a causal graph alone in the time-dependent setting.
@@
In this paper, we extend previous results by proposing an alternative definition of a sufficient time-dependent adjustment set and a novel criterion for comparing the variance of estimators based on different time-dependent adjustment sets. We show theoretically and empirically that employing our criterion can further decrease the variance of estimators of time-dependent treatment effects. We further show that our alternative definition of a sufficient time-dependent adjustment set allows for the identification of optimal time-dependent adjustment sets based on the graph alone. The rest of the paper is organized as follows. In Section 2 we review some background concepts, and present the assumed data structure and the notation. In Section 3 we provide a novel definition of time-dependent adjustment sets. Section 4 presents our theoretical results and numerical illustrations. Section 5 discusses possible applications of our results and potential extensions to our work. All proofs of results stated in the main text can be found in the Appendix.
@@ 2 Background
We begin by briefly reviewing some concepts of causal graph theory and introducing some notation. The unfamiliar reader is referred to other references for a more detailed introduction (e.g., [1] or the appendix of [6]). Let $\mathcal{G}$ be a directed acyclic graph (DAG) with nodes $\boldsymbol{V}$ representing the variables, that is, a graph for which all edges are unidirectional arrows and for which there is no cycle. A path is defined as a sequence of variables connected by arrows, independently of their direction. A causal path is a path in which all arrows point in the same direction. Opposingly, non-causal paths are paths in which not all arrows point in the same direction. A variable is said to be a collider on a given path if two arrows point to this variable on this path. Denote by $P$ the unknown distribution of the variables represented in the graph, and by $pa_{\mathcal{G}}(v_{k})$ the parents of $v_{k}$ according to the graph $\mathcal{G}$, that is, $pa_{\mathcal{G}}(v_{k})$ is the set of variables in $\mathcal{G}$ that have arrows stemming from them and pointing into $v_{k}$. Descendants of a node $v_{k}$ are all the nodes connected to $v_{k}$ by a causal path that starts with $v_{k}$. We say that a DAG is a causal graph if all common causes of the variables depicted in the graph are also on the graph [1] and if
$P=\prod_{v_{k}\in\boldsymbol{V}}f\left(v_{k}\mid pa_{\mathcal{G}}(v_{k})\right),$
@@
Suppose that we have a study with $T+1$ follow-up times and $n$ individuals sampled from a population so that the time-ordered (or otherwise topologically sorted) data take the following form $\mathcal{O}=(\boldsymbol{X}_{0},A_{0},\boldsymbol{X}_{1},A_{1},\dots,\boldsymbol{X}_{T},A_{T},Y)$ where $\boldsymbol{A}=(A_{0},...,A_{T})\subset\boldsymbol{V}$ is a time-varying exposure or treatment, $\boldsymbol{X}_{t}=(X_{t1},X_{t2},...,X_{tn_{xt}}),\ t=0,...,T$ is a vector of time-dependent covariates (with $n_{xt}$ the number of covariates at time $t$) and $Y\subset\boldsymbol{V}$ is the outcome of interest, measured at the end of the study. We assume that $A_{t},t=0,...,T$, correspond to finite valued random variables. For time ordered sets of variables, we use an overbar to denote the history of that variable up to a given time point. For example $\bar{\boldsymbol{A}}_{T}=(A_{0},...,A_{T})$. Let $Y^{\boldsymbol{a}}$ be the counterfactual outcome, that is, the outcome that would have been observed if, possibly contrary to fact, exposure had been $\boldsymbol{A}=\boldsymbol{a}$ where $\boldsymbol{a}=(a_{0},...,a_{T})$.
Based on a causal graph, it is possible to construct a single world intervention graph (SWIG) that represents the causal model (including counterfactual variables) under a hypothetical intervention [7]. As will be seen shortly, we will be using these SWIGs to identify sets of variables that meet the sequential conditional exchangeability assumption necessary to identify time-dependent causal effects. A SWIG can be constructed from a DAG by splitting each treatment node $A_{0},...,A_{T}$ in two parts, one part for the observed value and one for the value set by the (hypothetical) intervention: $A_{0}\mid a_{0},...,A_{T}\mid a_{T}$. All arrows entering into a node $A_{t}$ in the DAG also enter $A_{t}$ in the SWIG, whereas all arrows stemming from a node $A_{t}$ in the DAG stem from $a_{t}$ in the SWIG. Further, all descendants of a node $a_{t}$ in the SWIG are indexed by an $a_{t}$ superscript to emphasize that it is a counterfactual variable. Figure 1 presents an example of a DAG and its associated SWIG.
$X_{01}$$X_{02}$$A_{0}$$X_{11}$$X_{12}$$A_{1}$$Y$
@@
Independence between variables can be read from a causal DAG or a SWIG using the concept of d-separation [1]. Consider disjoint nodes, $V_{j}$ and $V_{k}$, and a set of nodes $\boldsymbol{V}_{\ell}$. The nodes $V_{j}$ and $V_{k}$ are d-separated by $\boldsymbol{V}_{\ell}$ under a graph $\mathcal{G}$ if all paths between $V_{j}$ and $V_{k}$ are blocked given $\boldsymbol{V}_{\ell}$ and we write $V_{j}\perp\!\!\!\perp_{\mathcal{G}}V_{k}\mid\boldsymbol{V}_{\ell}$. A path is said to be blocked given $\boldsymbol{V}_{\ell}$ if a) it contains a non-collider that is an element of $\boldsymbol{V}_{\ell}$, or b) there is a collider on the path where neither it nor its descendants are in $\boldsymbol{V}_{\ell}$. In the following, we adopt the convention that sets are d-separated from empty sets (i.e., $V_{j}\perp\!\!\!\perp_{\mathcal{G}}\emptyset$ and $V_{j}\perp\!\!\!\perp_{\mathcal{G}}\emptyset\mid\boldsymbol{V}_{\ell}$ are always true). If $V_{j}$ and $V_{k}$ are d-separated given $\boldsymbol{V}_{\ell}$, then $V_{j}$ and $V_{k}$ are also independent conditional on $\boldsymbol{V}_{\ell}$.
Our focus is on the efficient estimation of contrasts between counterfactual outcome expectations of the form $\Delta(P;\mathcal{G})\equiv\sum_{\boldsymbol{a}\in\mathcal{A}}c_{\boldsymbol{a}}\mathbb{E}_{P}(Y^{\boldsymbol{a}})$, where $\boldsymbol{a}$ is a vector taking values in the finite set $\mathcal{A}$ of all possible values that $\boldsymbol{A}$ can take and $c_{\boldsymbol{a}}\in\mathbb{R}$. For example, in the context of a two time point study with a binary treatment, $\boldsymbol{A}=(A_{0},A_{1})$ and $\mathcal{A}=\{(0,0),(1,1),(1,0),(0,1)\}$. If $c_{(0,0)}=-1$, $c_{(1,1)}=1$, and $c_{(1,0)}=c_{(0,1)}=0$ then $\Delta(P;\mathcal{G})=\mathbb{E}_{P}(Y^{(1,1)})-\mathbb{E}_{P}(Y^{(0,0)})$. These causal contrasts can be identified from the observed data by first identifying the counterfactual expectations $\chi_{\boldsymbol{a}}(P;\mathcal{G})\equiv\mathbb{E}_{P}(Y^{\boldsymbol{a}})$.
@@
Let $\boldsymbol{K}=(\boldsymbol{K}_{0},...,\boldsymbol{K}_{T})\subset\boldsymbol{X}$} be time ordered, disjoint sets of variables. Similar to Rotnitzky and Smucler [5], we say that $\boldsymbol{K}$ is a sufficient time-dependent adjustment set if
$\mathbb{E}_{P}(Y^{\boldsymbol{a}})$ $=\mathbb{E}_{P}(\mathbb{E}_{P}[...\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid\bar{\boldsymbol{A}}_{T}=\bar{\boldsymbol{a}}_{T},\bar{\boldsymbol{K}}_{T})\mid\bar{\boldsymbol{A}}_{T-1}=\bar{\boldsymbol{a}}_{T-1},\boldsymbol{\bar{K}}_{T-1}\}...\mid A_{0}=a_{0},\boldsymbol{K}_{0}])$ (1)
@@
$\mathbb{E}_{P}(Y^{\boldsymbol{a}})$ $=\mathbb{E}_{P}\left\{\frac{I_{\boldsymbol{a}}(\boldsymbol{A})Y}{\prod_{t=0}^{T}{P}(A_{t}=a_{t}\mid\bar{\boldsymbol{A}}_{t-1}=\bar{\boldsymbol{a}}_{t-1},\bar{\boldsymbol{K}}_{t})}\right\},$
@@
One way to identify a sufficient time-dependent adjustment set from a causal graph is to first construct a SWIG from the causal graph, then identify a set $\boldsymbol{K}_{0},...,\boldsymbol{K}_{T}$ such that the sequential conditional exchangeability assumption, $Y^{\boldsymbol{a}}\perp\!\!\!\perp A_{t}\mid\bar{\boldsymbol{A}}_{t-1},\bar{\boldsymbol{K}}_{t}$} for t=0,,Tt=0,...,T, holds, where $\bar{A}_{-1}\equiv\emptyset$. Example 1 illustrates this process.
@@
Assume that Figure 1 represents a causal DAG and its associated SWIG. Using d-separation rules, it can be seen that $Y^{\boldsymbol{a}}\perp\!\!\!\perp A_{0}\mid X_{02}$ and that $Y^{\boldsymbol{a}}\perp\!\!\!\perp A_{1}^{a_{0}}\mid X_{12}^{a_{0}},X_{02},A_{0}$, for example. Under consistency (i.e., if $A_{t}=a_{t}$ then $V^{a_{t}}=V$), the second conditional independence becomes $Y^{\boldsymbol{a}}\perp\!\!\!\perp A_{1}\mid X_{12},X_{02},A_{0}$}. As such, $\mathbb{E}_{P}(Y^{\boldsymbol{a}})=\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid A_{1}=a_{1},X_{12},A_{0}=a_{0},X_{02})\mid A_{0}=a_{0},X_{0,2}\}]$ and $\boldsymbol{K}=(\boldsymbol{K}_{0}=X_{02},\boldsymbol{K}_{1}=X_{12})$ is a time-dependent adjustment set.
@@ 3 Proposal
3.1 Extended definition of sufficient time-dependent adjustment sets
@@
Let $\boldsymbol{Z}=(\boldsymbol{Z}_{0},...,\boldsymbol{Z}_{T})$, with $\boldsymbol{Z}_{t}\subseteq\{\bar{\boldsymbol{X}}_{t},\bar{\boldsymbol{A}}_{t-1}\}$, t=0,,Tt=0,...,T. We say that $\boldsymbol{Z}$ is a sufficient time-dependent adjustment set if
$\mathbb{E}_{P}(Y^{\boldsymbol{a}})$ $=\mathbb{E}_{P}(\mathbb{E}_{P}[...\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid A_{t}=a_{T},\boldsymbol{Z}_{T}(\bar{\boldsymbol{a}}_{T-1}))\mid A_{T-1}=a_{T-1},\boldsymbol{Z}_{T-1}(\bar{\boldsymbol{a}}_{T-2})\}...\mid A_{0}=a_{0},\boldsymbol{Z}_{0}])$ (2)
@@
$\mathbb{E}_{P}(Y^{\boldsymbol{a}})$ $=\mathbb{E}_{P}\left\{\frac{I_{\boldsymbol{a}}(\boldsymbol{A})Y}{\prod_{t=0}^{T}{P}(A_{t}=a_{t}\mid\boldsymbol{Z}_{t}(\bar{\boldsymbol{a}}_{t-1}))}\right\}.$
@@
where we use the notation $\boldsymbol{Z}_{t}(\bar{\boldsymbol{a}}_{t-1})$ to emphasize that if some or all past treatments are in $\boldsymbol{Z}_{t}(\bar{\boldsymbol{a}}_{t-1})$ then they are held fixed.
@@
Definition 2 differs from Definition 1 in a few subtle but important aspects. First, unlike Definition 1, Definition 2 allows, but does not require, for previous treatments to be members of sufficient adjustment sets (i.e., $\bar{\boldsymbol{A}}_{t-1}\in\bar{\boldsymbol{Z}}_{t}$, t=1,,Tt=1,...,T, is allowed but not required in Definition 2, whereas $\bar{\boldsymbol{A}}_{t-1}$ is required to be conditioned on in Definition 1). In addition, the expectations in the identification formula (2) are only conditional on $A_{t}$ and separate $\boldsymbol{Z}_{t}$ at each time point, not on the histories $\bar{\boldsymbol{A}}_{t}$ and $\bar{\boldsymbol{K}}_{t}$, t=0,,Tt=0,...,T, as in Definition 1. Importantly, all sets that satisfy Definition 1 also satisfy Definition 2, but the reverse is false.
3.2 Construction of a sufficient time-dependent adjustment set according to Definition 2